. (2/5) ×5 whole 1 by 5  is equal to:
(a) 26/25 (b) 52/25 (c) 2/5 (d) 6

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Option (b) is correct here
First we have to convert the mixed fraction into improper fraction

5 whole 1/5 = 26/5
Then, (2/5) × (26/5)

= 52/25


Thanks